# InsurTech innovation using natural language processing

### v2

@@ 4.3 Standalone NLP Use Cases
In the actuarial science literature, there are relatively few standalone NLP use cases to date. For instance, [110] apply traditional text mining and classification techniques to customer service call transcripts, enabling insurers to categorize and process calls more efficiently—ultimately saving time and resources. [148] develop an NLP-powered repository and search engine focused on cyber risk literature, culminating in a web-based tool that facilitates access to relevant research in this emerging area. Additionally, [86] leverage legal documents to build a claim dispute prediction model using NLP, offering a novel approach to quantifying insurers’ litigation risk. [109] propose a LLM-powered framework to assist insurance claims processing by classifying insurance claims using distance-based metrics and analyzing semantic relationships through prompt-based methods. [77] explore potential actuarial applications of LLMs, including end-to-end claim management, report summarization, and drafting, through illustrative case studies emphasizing fraud detection and prevention. [105] propose an integrated and interactive framework that leverages Ethereum smart contracts for decentralized health insurance data storage and LLMs for detecting and analyzing insurance fraud in textual documents. [143] conduct three comprehensive case studies demonstrating that LLMs can serve as cost-efficient, effective, and nearly unbiased tools for automated claim handling, substantially reducing insurers’ operational overhead and streamlining workflows. Although limited in number, these studies demonstrate the versatility and practical value of NLP in insurance applications, suggesting ample opportunities for future exploration. To stimulate discussion and inspire future development, we present several representative standalone use cases of NLP in actuarial and insurance contexts. Given the rapid advancement in LLMs, this list is by no means exhaustive and is expected to grow over time.
@@
Claim Management: Claim management is a core function in insurance, encompassing reporting, investigation, review, and resolution. Traditionally, these processes have relied heavily on human expertise to ensure accuracy, largely due to the absence of effective automated tools capable of processing textual data (e.g., claim descriptions). Recent advancements in NLP, particularly the superior performance of LLMs on text-based tasks, have revealed the potential for automating claim management through LLM-powered systems [77, 143]. Moreover, [143] demonstrate that LLM-based claim management systems can achieve higher efficiency and cost-effectiveness, and in some cases, outperform human experts, indicating substantial potential for reducing operational overhead and streamlining insurance workflows.
Fraud Detection and Prevention: Fraud has long been a persistent concern in financial sectors, particularly in the insurance industry. Although inherently difficult to detect and analyze, traditional analytical approaches grounded in economic theory, game-theoretic modeling, and incentive design—as developed by [84, 83, 115, 120]—seek to explain the underlying mechanisms of fraudulent behavior and propose strategies to mitigate issues such as information asymmetry, planned fraud, moral hazard, and adverse selection. With the growth of data availability, data-driven methods have increasingly been employed for fraud detection and prevention. In particular, the recent advancement of LLMs has enabled the analysis of subtle fraudulent patterns embedded in textual data, which provide richer contextual and sentiment information for more accurate data-driven fraud identification [77, 105].
@@ 4.3.1 Business Industry Classification
One of the most critical rating factors in commercial insurance pricing is industry classification, i.e., the industry to which a business belongs. This classification serves as the foundation for underwriting and risk assessment, as risk exposure varies significantly across industries. While the process might appear straightforward, accurately classifying a business’s industry is far more complex than it seems. Currently, insurers primarily rely on self-reported information from business owners [94], combined with underwriters’ professional judgment, to determine industry classifications. However, this approach often lacks consistency and structure, leading to potential misclassifications. Such subjectivity not only introduces pricing unfairness but also exposes insurers to unforeseen risks due to incorrect industry classification. Adding to the challenge is the fact that many modern businesses span multiple industries. For example, consider a brewery: it manufactures alcoholic beverages (suggesting a manufacturing classification), sells bottled products directly to consumers (fitting a retail profile), and may also operate an on-site bar or restaurant (aligning with food services). Determining a single industry classification for such a multifaceted operation is not only difficult for business owners but also confusing for underwriters. Misclassification in these cases can lead to inappropriate pricing, coverage gaps, or even regulatory issues. As businesses become increasingly hybrid and diversified, the need for a more systematic, data-driven approach to industry classification becomes crucial to ensuring both actuarial fairness and underwriting accuracy.
@@
From an NLP technical perspective, there are two main approaches to building an automated business industry classification system. The first approach is to frame the problem as a text classification task, where given a piece of text, the model predicts the group or class it belongs to. In fact, sentiment analysis, which we discussed earlier in Section 4.2.2, is a specific form of text classification. Even language modeling mentioned in Section 3.2 can be viewed through this lens: predicting the next word is essentially classifying the current context into one of many possible word classes. A typical text classification pipeline involves three major steps: feature extraction, dimensionality reduction (if necessary), and classification. However, this approach requires a well-labeled dataset where each piece of text is clearly assigned to a target class. For example, the BEACON system, an industry classification tool proposed by [94], is trained on proprietary U.S. Census data containing 4.3 million observations, which enables it to adopt a supervised learning approach. In our case, the dataset contains information for $13,873$ businesses. Among them, only $2,372$ of these businesses have confirmed NAICS codes and proper descriptive information available on their official webpages. If we were to pursue a supervised text classification approach, we would have to discard the majority of our data due to missing labels or non-informative features, significantly limiting the amount of training data available. Furthermore, the vast number of categories at the four- and six-digit levels results in severe underrepresentation for most of the codes, raising concerns about the reliability and credibility of the resulting model.
@@
Therefore, we adopt the second approach: unsupervised learning leveraging topic modeling. Topic modeling aims to uncover latent thematic structures within large collections of text by analyzing the co-occurrence patterns of words across documents. Through this process, it identifies coherent sets of words that together represent distinct “topics.” In our study, we employ well-known and interpretable topic modeling techniques: Latent Dirichlet Allocation (LDA) [80] and Rapid Automatic Keyword Extraction (RAKE) [128]. LDA is a generative probabilistic model designed to represent documents as mixtures of topics, where each topic is characterized by a distribution over words. Specifically, LDA assumes that each document $d_{j}$, $j=1,\dots,n$ is a mixture of a finite number of topics $z_{k}$, $k=1\dots Z$, and each topic is characterized by a distribution over words. The main parameters of LDA include $\theta_{j}$ and $\phi_{k}$. $\theta_{j}$ is the topic distribution for document $d_{j}$, $j=1,\dots,n$ and it is drawn from a Dirichlet distribution with parameter $\alpha$, i.e., $\theta_{j}\sim\text{Dirichlet}(\alpha)$. $\phi_{k}$ represents the word distribution for topic $z_{k}$, $k=1\dots Z$, drawn from a Dirichlet distribution with parameter $\beta$, i.e., $\phi_{k}\sim\text{Dirichlet}(\beta)$. For each word $t_{i}$, $i=1,\dots,m$ in the document $d_{j}$, $j=1,\dots,n$, a topic $z_{d_{j},t_{i}}$ is chosen from a multinomial distribution $\text{Multinomial}(\theta_{j})$, and a word $t_{i}$ is selected based on $\text{Multinomial}(\phi_{z_{d_{j},t_{i}}})$. Mathematically, the joint distribution over topics, words, and documents in LDA is given by:
@@
Using the embedding mentioned in Section 3.2, we represent each business as vectors based on its InsurTech categories and extracted topic words. A similar embedding process is applied to the NAICS code descriptions. While various embedding choices are available, we omit detailed comparisons here and refer interested readers to [86] for further discussion. We select the highest-dimensional version of the chosen embedding model, as higher-dimensional vectors typically capture more semantic information and yield better performance in downstream tasks. Specifically, we adopt the MiniLM embedding model [140] provided by sentence_transformer [^8] https://sbert.net/#.
@@
As one of the comparable benchmark, BEACON reports a Top-1 accuracy of 0.415 and a Top-3 accuracy of 0.614 for 2-digit NAICS code classification (Ensemble, Table 4 in [94]). In contrast, our unsupervised learning approach achieves broadly comparable or even better performance while using far less data and operating more efficiently. It is also worth emphasizing that BEACON relies on a much larger training set and an entirely different methodology (supervised learning), yet the results for the industry classification are of a similar order of magnitude. Furthermore, motivated by the recent surge of LLMs as the first-choice NLP solution across various domains, we also investigate their potential for industry classification and consider it as another comparable benchmark. A comparison of its performance against our unsupervised learning approach is presented in Table 8. Our experiments reveal that general-purpose LLMs, without task-specific fine-tuning or modification, struggle even to correctly interpret and execute the industry classification task. In contrast, our unsupervised learning solution remains the more reliable choice: it is substantially more computationally efficient and provides a practical and interpretable decision-support system with measurable prediction probabilities. For a detailed description of the locally deployed open-source LLM experiments using prompt engineering and Retrieval-Augmented Generation (RAG) enhanced LLM, please refer to Appendix D. Due to data privacy, we opt out of using API calls from various commercial LLMs.
@@ Appendix D Setting of LLM Industry Classification
The rapid rise of LLM-powered Generative AI (GenAI) systems has enabled a wide range of applications, particularly in text generation. A natural question is whether LLMs can also assist in industry classification and how their performance compares to our unsupervised learning solution. To investigate this, we design a series of experiments on LLM-based industry classification. Due to the sensitivity of our dataset, we restrict our evaluation to open-source models rather than calling commercial closed-source APIs. In particular, we adopt the 32B variant of DeepSeek-R1 [99], one of the strongest performing open-source LLMs, to generate classifications based on firm categorizations and descriptions.

### v3

@@ 4.3 Standalone NLP Use Cases
In the actuarial science literature, there are relatively few standalone NLP use cases to date. For instance, [Liao2020] apply traditional text mining and classification techniques to customer service call transcripts, enabling insurers to categorize and process calls more efficiently—ultimately saving time and resources. [Zhang11112024] develop an NLP-powered repository and search engine focused on cyber risk literature, culminating in a web-based tool that facilitates access to relevant research in this emerging area. Additionally, [cao2024assessing] leverage legal documents to build a claim dispute prediction model using NLP, offering a novel approach to quantifying insurers’ litigation risk. [liTextualAnalysisInsurance2025] propose a LLM-powered framework to assist insurance claims processing by classifying insurance claims using distance-based metrics and analyzing semantic relationships through prompt-based methods. Although limited in number, these studies demonstrate the versatility and practical value of NLP in insurance applications, suggesting ample opportunities for future exploration. To stimulate discussion and inspire future development, we present several representative standalone use cases of NLP in actuarial and insurance contexts. Given the rapid advancement in LLMs, this list is by no means exhaustive and is expected to grow over time.
@@ 4.3.1 Business Industry Classification
One of the most critical rating factors in commercial insurance pricing is industry classification, i.e., the industry to which a business belongs. This classification serves as the foundation for underwriting and risk assessment, as risk exposure varies significantly across industries. While the process might appear straightforward, accurately classifying a business’s industry is far more complex than it seems. Currently, insurers primarily rely on self-reported information from business owners [dumbacherBEACONToolIndustry2025], combined with underwriters’ professional judgment, to determine industry classifications. However, this approach often lacks consistency and structure, leading to potential misclassifications. Such subjectivity not only introduces pricing unfairness but also exposes insurers to unforeseen risks due to incorrect industry classification. Adding to the challenge is the fact that many modern businesses span multiple industries. For example, consider a brewery: it manufactures alcoholic beverages (suggesting a manufacturing classification), sells bottled products directly to consumers (fitting a retail profile), and may also operate an on-site bar or restaurant (aligning with food services). Determining a single industry classification for such a multifaceted operation is not only difficult for business owners but also confusing for underwriters. Misclassification in these cases can lead to inappropriate pricing, coverage gaps, or even regulatory issues. As businesses become increasingly hybrid and diversified, the need for a more systematic, data-driven approach to industry classification becomes crucial to ensuring both actuarial fairness and underwriting accuracy.
@@
From an NLP technical perspective, there are two main approaches to building an automated business industry classification system. The first approach is to frame the problem as a text classification task, where given a piece of text, the model predicts the group or class it belongs to. In fact, sentiment analysis, which we discussed earlier in Section 4.2.2, is a specific form of text classification. Even language modeling mentioned in Section 3.2 can be viewed through this lens: predicting the next word is essentially classifying the current context into one of many possible word classes. A typical text classification pipeline involves three major steps: feature extraction, dimensionality reduction (if necessary), and classification. However, this approach requires a well-labeled dataset where each piece of text is clearly assigned to a target class. For example, the BEACON system, an industry classification tool proposed by [dumbacherBEACONToolIndustry2025], is trained on proprietary U.S. Census data containing 4.3 million observations, which enables it to adopt a supervised learning approach. In our case, the dataset contains information for $13,873$ businesses. Among them, only $2,372$ of these businesses have confirmed NAICS codes and proper descriptive information available on their official webpages. If we were to pursue a supervised text classification approach, we would have to discard the majority of our data due to missing labels or non-informative features, significantly limiting the amount of training data available. Furthermore, the vast number of categories at the four- and six-digit levels results in severe underrepresentation for most of the codes, raising concerns about the reliability and credibility of the resulting model.
@@
Therefore, we adopt the second approach: unsupervised learning leveraging topic modeling. Topic modeling aims to uncover latent thematic structures within large collections of text by analyzing the co-occurrence patterns of words across documents. Through this process, it identifies coherent sets of words that together represent distinct “topics.” In our study, we employ well-known and interpretable topic modeling techniques: Latent Dirichlet Allocation (LDA) [Blei2003] and Rapid Automatic Keyword Extraction (RAKE) [Stuart2010]. LDA is a generative probabilistic model designed to represent documents as mixtures of topics, where each topic is characterized by a distribution over words. Specifically, LDA assumes that each document $d_{j}$, $j=1,\dots,n$ is a mixture of a finite number of topics $z_{k}$, $k=1\dots Z$, and each topic is characterized by a distribution over words. The main parameters of LDA include $\theta_{j}$ and $\phi_{k}$. $\theta_{j}$ is the topic distribution for document $d_{j}$, $j=1,\dots,n$ and it is drawn from a Dirichlet distribution with parameter $\alpha$, i.e., $\theta_{j}\sim\text{Dirichlet}(\alpha)$. $\phi_{k}$ represents the word distribution for topic $z_{k}$, $k=1\dots Z$, drawn from a Dirichlet distribution with parameter $\beta$, i.e., $\phi_{k}\sim\text{Dirichlet}(\beta)$. For each word $t_{i}$, $i=1,\dots,m$ in the document $d_{j}$, $j=1,\dots,n$, a topic $z_{d_{j},t_{i}}$ is chosen from a multinomial distribution $\text{Multinomial}(\theta_{j})$, and a word $t_{i}$ is selected based on $\text{Multinomial}(\phi_{z_{d_{j},t_{i}}})$. Mathematically, the joint distribution over topics, words, and documents in LDA is given by:
@@
Using the embedding mentioned in Section 3.2, we represent each business as vectors based on its InsurTech categories and extracted topic words. A similar embedding process is applied to the NAICS code descriptions. While various embedding choices are available, we omit detailed comparisons here and refer interested readers to [cao2024assessing] for further discussion. We select the highest-dimensional version of the chosen embedding model, as higher-dimensional vectors typically capture more semantic information and yield better performance in downstream tasks. Specifically, we adopt the MiniLM embedding model [wangMINILMDeepSelfattention2020] provided by sentence_transformer [^8] https://sbert.net/#.
@@
As one of the comparable benchmark, BEACON reports a Top-1 accuracy of 0.415 and a Top-3 accuracy of 0.614 for 2-digit NAICS code classification (Ensemble, Table 4 in [dumbacherBEACONToolIndustry2025]). In contrast, our unsupervised learning approach achieves broadly comparable or even better performance while using far less data and operating more efficiently. It is also worth emphasizing that BEACON relies on a much larger training set and an entirely different methodology (supervised learning), yet the results for the industry classification are of a similar order of magnitude. Furthermore, motivated by the recent surge of LLMs as the first-choice NLP solution across various domains, we also investigate their potential for industry classification and consider it as another comparable benchmark. A comparison of its performance against our unsupervised learning approach is presented in Table 8. Our experiments reveal that general-purpose LLMs, without task-specific fine-tuning or modification, struggle even to correctly interpret and execute the industry classification task. In contrast, our unsupervised learning solution remains the more reliable choice: it is substantially more computationally efficient and provides a practical and interpretable decision-support system with measurable prediction probabilities. For a detailed description of the locally deployed open-source LLM experiments using prompt engineering and Retrieval-Augmented Generation (RAG) enhanced LLM, please refer to Appendix D. Due to data privacy, we opt out of using API calls from various commercial LLMs.
@@ Appendix D Setting of LLM Industry Classification
The rapid rise of LLM-powered Generative AI (GenAI) systems has enabled a wide range of applications, particularly in text generation. A natural question is whether LLMs can also assist in industry classification and how their performance compares to our unsupervised learning solution. To investigate this, we design a series of experiments on LLM-based industry classification. Due to the sensitivity of our dataset, we restrict our evaluation to open-source models rather than calling commercial closed-source APIs. In particular, we adopt the 32B variant of DeepSeek-R1 [guoDeepSeekR1IncentivizesReasoning2025], one of the strongest performing open-source LLMs, to generate classifications based on firm categorizations and descriptions.